OpenStudy (anonymous):

polar coordinates of (-1, -2) and (1, 3) that satisfy 0≤theta≤2pi and -pi ≤theta≤pi ?

OpenStudy (anonymous):

use the formula..\[r=\sqrt{x^2+y^2}\] and \[\theta =\tan^{-1} (\frac yx)\]

OpenStudy (anonymous):

idk if i got the answers right...

OpenStudy (anonymous):

1st prob. r=sqrt(5) theta=pi+tan^-1(2) first case and theta= -tan^-1(2) for second case...